A 15 gram rubber balloon is filled with 18 L of Helium (density 0.179 kg/m3) and tied with a string to a child's wrist. Find the tension in the cord.


1
Expert's answer
2020-09-16T10:10:36-0400


The free body diagram is shown on the figure. Here:

"\mathbf{F}_b = \rho_{air}\cdot g\cdot V_{He}" is the buoyant force, "T" is the tension in the cord and "\mathbf{mg}" is the gravity force.

In projections on the vertical axis get:

"F_b - mg - T = 0\\
T = F_b - mg = \rho_{air}\cdot g\cdot V_{He} - mg =g ( \rho_{air}\cdot V_{He} - m)"

Let's substitute the numerical values: "\rho_{air} = 1.2 kg/m^3", "V_{He} = 18L = 0.018 m^3" and "g = 9.81m\/s^2".

Here "m" is the mass of the ballon together with 18 L of Helium:


"m = 0.015kg + V_{He}\cdot \rho_{He} = 0.015+ 0.018\cdot 0.179 = 0.018222\space kg"

Finally, obtain:


"T =g ( \rho_{air}\cdot V_{He} - m) =9.81\cdot ( 1.2\cdot 0.018 - 0.018222) \approx 0.033N"

Answer. 0.033 N.